Net Income Planning Nolden Company has charged a selling price of $20 per unit, incurred variable costs of $12 per unit, and total fixed costs of $95,000. What unit sales volume is necessary to earn the following related amounts of net income before income tax? (a) $18,000; (b) $27,000, or (c) equal to 20% of sales revenue. (a) 0 (b) 0 (c) 0 units units units Check
Answer #1
Sales volume (unit)                  =( fixed cost+ target income)/contribution margin per unit
a) (95000+18000)/(20-12)
14125 units answer
b) (95000+27000)/8
15250 units answer
c) let sales units be x
contribution = fixed cost + net income
8x                   = 95000 + 20x*20%
8x                    = 95000 + 4x
x                      = 95000/4
23750 units answer